# Digging into the Interior of Hot Cores with ALMA. VI. The Formation of Low-mass Multiple Systems in High-mass Cluster-forming Regions

### v1

@@ IV.3 Condensation Surface Density Effects
Multiplicity variations in nearby low-mass star-forming regions may largely reflect differences in stellar surface densities, which in turn could promote multiple system formation and lead to a higher MF (J. J. Tobin et al., 2022; S. S. R. Offner et al., 2023). We investigate this effect in our sample of condensations embedded in high-mass cluster-forming regions and find no correlation. Figure 4 (also Table LABEL:tab2) presents the condensation MF as a function of surface density, ranging from 40 to 433 pc-2. For comparison, the average surface density or condensations, as measured in Perseus and Orion, is 30 pc-2 (J. J. Tobin et al., 2022), whereas the average surface density for condensations in this study of high-mass regions is 210 pc-2. Figure 4 illustrates how the MF remains relatively constant across different condensation surface densities, regardless of whether all condensations are included or the sample is divided into those embedded in cores and orphans. The discrepancy between low-mass and high-mass regions may suggest that the condensation surface density plays a different role in shaping multiplicity. We propose that the more chaotic and dynamical conditions prevalent in high-mass environments disrupt the tendency observed in low-mass regions. At sufficiently high surface densities, these conditions may hinder the ability to retain binary systems, particularly those that are less gravitationally bound. This interpretation aligns with the cluster-forming simulations conducted by (D. Guszejnov et al., 2023), which show that multiplicity decreases with increasing local stellar density. Indeed, the discovery of the orphan population provides further evidence for the highly dynamical conditions in high-mass cluster-forming regions. Approximately 47% of the orphans are co-located with diffuse continuum emission, but are not within a core. This subpopulation may have been ejected from cores due to dynamical interactions. On the other hand, the remaining 53% of the orphan population is unassociated with diffuse dust continuum and appears to be more isolated. This subpopulation of orphans may be an older generation of protostars formed in low-mass cores that have already dispersed most of their envelope. As shown in Figure 6 and detailed in the Appendix C, as a proxy for the condensation mass in the optically thin regime, the millimeter luminosity of the orphan population is generally lower than that of condensations embedded in cores. With both, a lower mass distribution and a lower MF of 8%, the orphan population has properties consistent with the lowest-mass stars (0.1 M⊙) in the Main Sequence phase (S. S. R. Offner et al., 2023).
@@ A.1 Core Identification
Using Astrodendro 111https://dendrograms.readthedocs.io/en/stable/ Package, K. Ishihara et al. (2024) identify compact structures in the 1.33 mm dust continuum emission using the C43-5 configuration data (03). In their work, the ‘leaf’, the highest structure in the hierarchy tree, is defined as the ‘core’. The parameters obtained from K. Ishihara et al. (2024) include coordinates, flux density, core radius, and core area for all identified structures, as provided by Astrodendro. These results are summarized in the Table LABEL:tab3. A total of 415 cores are identified among the 23 regions studied, containing a mixed population that includes high-mass cores (including hot cores), and low-mass cores.
Fixed-size ‘cores’: To facilitate a comparison with nearby cloud statistics, where the core size remains relatively constant, we also adopt a uniform core size. A constant core size also allows us to assess the effect in the analysis of having cores with different sizes. However, the maximum size that could be adopted is restricted by the proximity of cores in the clustered environments under study. If the adopted size is excessively large, we will combine independent cores with their respective multiple systems. To explore what fixed core size could be used, we run the algorithm defining multiples up to 6000 au, ignoring core boundaries. Figure 5 presents a comparison of the condensation companion separation distribution (obtained from Appendix F) and the core separation distribution (obtained from K. Ishihara et al. (2024)), revealing an overlap between the two distributions. The results are only used to test whether the multiplicity statistics are sensitive to the choice of core size, and they are shown only in Figure 5. To mitigate contamination in the definition of multiple systems caused by condensations from nearby cores, we apply a fixed core size of 3000 au to all cores in the subsequent multiplicity analysis. The fixed size of 3000 au reduces the contamination from neighboring cores to 2.
@@ B.1 Core Mass
Here, Sν represents the flux density of the cores, is the distance to each region, is the gas-to-dust mass ratio of 100, and the dust opacity =0.899 cm2 g-1 corresponds to the opacity of dust grains with thin ice mantles at a gas density of cm-3 (V. Ossenkopf & T. Henning, 1994). is is the Planck function for a blackbody at the dust temperature . We adopt the hot core temperature determined by (K. Taniguchi et al., 2023) as the dust temperatures when available. For the remaining cores within a given region, we adopt the single-dish dust temperature reported by K. Ishihara et al. (2024). The distance and dust temperature information in Table LABEL:tab2.
The masses of the identified cores range from 0.071 to 33.5 M⊙. To separate the core sample between high-mass and low-mass, we first assumed that all hot cores identified by K. Taniguchi et al. (2023) will likely form a high-mass star, independent of their masses. Cores with masses exceeding 16 M⊙ are classified as high-mass cores. The remaining ones, with masses below 16 M⊙ (excluding hot cores), are classified as low-mass cores. These cores have a mass ranging from 0.071 to 13.97 M⊙. As a result, we have 378 low-mass cores and 37 high-mass cores in the 23 regions.
@@ B.2 H2 Number Density of Cores
Here, represents the core mass, and is the core radius. = 2.8 (J. Kauffmann et al., 2008) is the mean molecular weight, and is the mass of a hydrogen atom. The number density of cores is calculated and grouped into two categories: one containing cores that formed single systems, and another one containing cores that formed multiple systems.
@@
We note that the uncertainty in the mass and the density is dominated by the uncertainty of the gas-to-dust mass ratio and the dust opacity, with 23% and 28% being the respective values (P. Sanhueza et al., 2017, 2019), resulting in a mass and number density uncertainty of about 50% their value.
@@ Appendix E Measuring Multiplicity
To gather the multiplicity statistics of low-mass condensations in high-mass cluster-forming regions, we adopt the formulation from (J. J. Tobin et al., 2022) to calculate two key parameters, the multiplicity fraction (MF) and the companion fraction (CF), along with their uncertainties, as shown below:
@@ Appendix F Calculating Companion Separation
To analyze the separation between member condensations in a multiple system, we adopt the inside-out search method from J. J. Tobin et al. (2022) for a consistent comparison with previous observations and simulations. The procedure for calculating the companion separations in each cluster-forming region is as follows: (1) The search begins from any condensation within a core, with a minimum search radius of 50 au. (2) We search for companions for each condensation by incrementally increasing the radius in 10 au steps, extending the search up to a defined boundary.
@@
To examine the effects of the core definition, we search for multiples ignoring core’s boundaries. We calculate the companion separation blindly to all condensations up to 6000 au. Figure 5 compares the resulting companion separation distribution with the core separation distribution derived by K. Ishihara et al. (2024) using the minimum spanning tree (MST) algorithm. The MST algorithm connects a set of core centers using the shortest possible network of branches without forming loops. The companion separation distribution reveals two peaks: one at 103.1 au (1200 au) and another around 103.6 au(4000 au). The second peak coincides with the peak of the core separation distribution from (K. Ishihara et al., 2024), indicating that if multiples are defined blindly, without a boundary, condensations embedded in different cores will be considered as multiples. Therefore, we interpret the 1200 au peak as a robust characteristic fragmentation scale and the true companion separation.
@@ Appendix H Measuring Chance Alignment
To assess the reliability of companion separations and exclude the impact of random alignment along the line of sight, we employ the method from J. J. Tobin et al. (2022) to calculate the probability of being a real companion (the complement of the probability of chance alignment) for each pair of condensation companions.
@@
Here, = 0.4 accounts for the observational completeness, which we obtained by comparing the condensation flux density in our sample with that from J. J. Tobin et al. (2022). In their work, the completeness was estimated as 75% based on Lupus surveys (M. Ansdell et al., 2016), but applying the same approach to our sample yields a lower value of 40%. The variable represents the separation between two condensation companions, measured in . For the surface density, we assume as a representative field stellar surface density the one from the ONC, = 770 pc-2 (B. Reipurth et al., 2007). Additionally, in Figure 7, we adopt the condensation surface density derived for each specific field (Appendix G). This companion separation distribution exhibits a similar shape and also peaks around 1000 au.
@@ Appendix I Initial Fragmentation Scale
We apply the approach described in R. L. Kuruwita & T. Haugbølle (2023) to estimate the initial fragmentation scale as a proxy for turbulent core fragmentation, which approximates how the star-forming environment influences the separation of multiple systems within cores considering the surrounding density and turbulence. This approach assumes that cores, on average, can be modeled as critical-mass Bonnor-Ebert spheres. The fragmentation scale is then approximated to the radius of a critical Bonnor-Ebert core modified by an efficiency factor, , as
in which is an efficiency factor that we consider to be between 0.55 (from simulations in R. L. Kuruwita & T. Haugbølle, 2023) and 1.0 (fragmentation of the core occurring at the edge of the core). The sound speed, , is given by
@@
in which is the clump density, and is the Mach number given by . is the non-thermal velocity dispersion derived from = and = is the observed velocity dispersion, where FWHM is the full width at half maximum of HN13C lines. The thermal velocity dispersion, , of HN13C molecules is calculated using Equation I2 by replacing for the molecular weight of HN13C, 28. We extracted the average line FWHM of the envelope surrounding each core using the optically thin HN13C transition line from the MagMaR (Magnetic Fields in Massive Star-forming Regions) project (M. Fernández-López et al., 2021; P. Sanhueza et al., 2021, 2025; P. C. Cortés et al., 2021, 2024; P. Saha et al., 2024; L. A. Zapata et al., 2024), which observed the same sample of this study at the same angular resolution than the C-5 configuration (03).

### v2

@@ IV.3 Condensation Surface Density Effects
Multiplicity variations in nearby low-mass star-forming regions may largely reflect differences in stellar surface densities, which in turn could promote multiple system formation and lead to a higher MF (Tobin et al., 2022; Offner et al., 2023). We investigate this effect in our sample of condensations embedded in high-mass cluster-forming regions and find no correlation. Figure 4 (also Table LABEL:tab2) presents the condensation MF as a function of surface density, ranging from 40 to 433 pc-2. For comparison, the average surface density or condensations, as measured in Perseus and Orion, is 30 pc-2 (Tobin et al., 2022), whereas the average surface density for condensations in this study of high-mass regions is 210 pc-2. Figure 4 illustrates how the MF remains relatively constant across different condensation surface densities, regardless of whether all condensations are included or the sample is divided into those embedded in cores and orphans. The discrepancy between low-mass and high-mass regions may suggest that the condensation surface density plays a different role in shaping multiplicity. We propose that the more chaotic and dynamical conditions prevalent in high-mass environments disrupt the tendency observed in low-mass regions. At sufficiently high surface densities, these conditions may hinder the ability to retain binary systems, particularly those that are less gravitationally bound. This interpretation aligns with the cluster-forming simulations conducted by (Guszejnov et al., 2023), which show that multiplicity decreases with increasing local stellar density. Indeed, the discovery of the orphan population provides further evidence for the highly dynamical conditions in high-mass cluster-forming regions. Approximately 47% of the orphans are co-located with diffuse continuum emission, but are not within a core. This subpopulation may have been ejected from cores due to dynamical interactions. On the other hand, the remaining 53% of the orphan population is unassociated with diffuse dust continuum and appears to be more isolated. This subpopulation of orphans may be an older generation of protostars formed in low-mass cores that have already dispersed most of their envelope. As shown in Figure 6 and detailed in the Appendix C, as a proxy for the condensation mass in the optically thin regime, the millimeter luminosity of the orphan population is generally lower than that of condensations embedded in cores. With both, a lower mass distribution and a lower MF of 8%, the orphan population has properties consistent with the lowest-mass stars (0.1 M⊙) in the Main Sequence phase (Offner et al., 2023).
@@ A.1 Core Identification
Using Astrodendro 111https://dendrograms.readthedocs.io/en/stable/ Package, Ishihara et al. (2024) identify compact structures in the 1.33 mm dust continuum emission using the C43-5 configuration data (03). In their work, the ‘leaf’, the highest structure in the hierarchy tree, is defined as the ‘core’. The parameters obtained from Ishihara et al. (2024) include coordinates, flux density, core radius, and core area for all identified structures, as provided by Astrodendro. These results are summarized in the Table LABEL:tab3. A total of 415 cores are identified among the 23 regions studied, containing a mixed population that includes high-mass cores (including hot cores), and low-mass cores.
Fixed-size ‘cores’: To facilitate a comparison with nearby cloud statistics, where the core size remains relatively constant, we also adopt a uniform core size. A constant core size also allows us to assess the effect in the analysis of having cores with different sizes. However, the maximum size that could be adopted is restricted by the proximity of cores in the clustered environments under study. If the adopted size is excessively large, we will combine independent cores with their respective multiple systems. To explore what fixed core size could be used, we run the algorithm defining multiples up to 6000 au, ignoring core boundaries. Figure 5 presents a comparison of the condensation companion separation distribution (obtained from Appendix F) and the core separation distribution (obtained from Ishihara et al. (2024)), revealing an overlap between the two distributions. The results are only used to test whether the multiplicity statistics are sensitive to the choice of core size, and they are shown only in Figure 5. To mitigate contamination in the definition of multiple systems caused by condensations from nearby cores, we apply a fixed core size of 3000 au to all cores in the subsequent multiplicity analysis. The fixed size of 3000 au reduces the contamination from neighboring cores to 2.
@@ B.1 Core Mass
Here, Sν represents the flux density of the cores, is the distance to each region, is the gas-to-dust mass ratio of 100, and the dust opacity =0.899 cm2 g-1 corresponds to the opacity of dust grains with thin ice mantles at a gas density of cm-3 (Ossenkopf and Henning, 1994). is is the Planck function for a blackbody at the dust temperature . We adopt the hot core temperature determined by (Taniguchi et al., 2023) as the dust temperatures when available. For the remaining cores within a given region, we adopt the single-dish dust temperature reported by Ishihara et al. (2024). The distance and dust temperature information in Table LABEL:tab2.
The masses of the identified cores range from 0.071 to 33.5 M⊙. To separate the core sample between high-mass and low-mass, we first assumed that all hot cores identified by Taniguchi et al. (2023) will likely form a high-mass star, independent of their masses. Cores with masses exceeding 16 M⊙ are classified as high-mass cores. The remaining ones, with masses below 16 M⊙ (excluding hot cores), are classified as low-mass cores. These cores have a mass ranging from 0.071 to 13.97 M⊙. As a result, we have 378 low-mass cores and 37 high-mass cores in the 23 regions.
@@ B.2 H2 Number Density of Cores
Here, represents the core mass, and is the core radius. = 2.8 (Kauffmann et al., 2008) is the mean molecular weight, and is the mass of a hydrogen atom. The number density of cores is calculated and grouped into two categories: one containing cores that formed single systems, and another one containing cores that formed multiple systems.
@@
We note that the uncertainty in the mass and the density is dominated by the uncertainty of the gas-to-dust mass ratio and the dust opacity, with 23% and 28% being the respective values (Sanhueza et al., 2017, 2019), resulting in a mass and number density uncertainty of about 50% their value.
@@ Appendix E Measuring Multiplicity
To gather the multiplicity statistics of low-mass condensations in high-mass cluster-forming regions, we adopt the formulation from (Tobin et al., 2022) to calculate two key parameters, the multiplicity fraction (MF) and the companion fraction (CF), along with their uncertainties, as shown below:
@@ Appendix F Calculating Companion Separation
To analyze the separation between member condensations in a multiple system, we adopt the inside-out search method from Tobin et al. (2022) for a consistent comparison with previous observations and simulations. The procedure for calculating the companion separations in each cluster-forming region is as follows: (1) The search begins from any condensation within a core, with a minimum search radius of 50 au. (2) We search for companions for each condensation by incrementally increasing the radius in 10 au steps, extending the search up to a defined boundary.
@@
To examine the effects of the core definition, we search for multiples ignoring core’s boundaries. We calculate the companion separation blindly to all condensations up to 6000 au. Figure 5 compares the resulting companion separation distribution with the core separation distribution derived by Ishihara et al. (2024) using the minimum spanning tree (MST) algorithm. The MST algorithm connects a set of core centers using the shortest possible network of branches without forming loops. The companion separation distribution reveals two peaks: one at 103.1 au (1200 au) and another around 103.6 au(4000 au). The second peak coincides with the peak of the core separation distribution from (Ishihara et al., 2024), indicating that if multiples are defined blindly, without a boundary, condensations embedded in different cores will be considered as multiples. Therefore, we interpret the 1200 au peak as a robust characteristic fragmentation scale and the true companion separation.
@@ Appendix H Measuring Chance Alignment
To assess the reliability of companion separations and exclude the impact of random alignment along the line of sight, we employ the method from Tobin et al. (2022) to calculate the probability of being a real companion (the complement of the probability of chance alignment) for each pair of condensation companions.
@@
Here, = 0.4 accounts for the observational completeness, which we obtained by comparing the condensation flux density in our sample with that from Tobin et al. (2022). In their work, the completeness was estimated as 75% based on Lupus surveys (Ansdell et al., 2016), but applying the same approach to our sample yields a lower value of 40%. The variable represents the separation between two condensation companions, measured in . For the surface density, we assume as a representative field stellar surface density the one from the ONC, = 770 pc-2 (Reipurth et al., 2007). Additionally, in Figure 7, we adopt the condensation surface density derived for each specific field (Appendix G). This companion separation distribution exhibits a similar shape and also peaks around 1000 au.
@@ Appendix I Initial Fragmentation Scale
We apply the approach described in Kuruwita and Haugbølle (2023) to estimate the initial fragmentation scale as a proxy for turbulent core fragmentation, which approximates how the star-forming environment influences the separation of multiple systems within cores considering the surrounding density and turbulence. This approach assumes that cores, on average, can be modeled as critical-mass Bonnor-Ebert spheres. The fragmentation scale is then approximated to the radius of a critical Bonnor-Ebert core modified by an efficiency factor, , as
in which is an efficiency factor that we consider to be between 0.55 (from simulations in Kuruwita and Haugbølle, 2023) and 1.0 (fragmentation of the core occurring at the edge of the core). The sound speed, , is given by
@@
in which is the clump density, and is the Mach number given by . is the non-thermal velocity dispersion derived from = and = is the observed velocity dispersion, where FWHM is the full width at half maximum of HN13C lines. The thermal velocity dispersion, , of HN13C molecules is calculated using Equation I2 by replacing for the molecular weight of HN13C, 28. We extracted the average line FWHM of the envelope surrounding each core using the optically thin HN13C transition line from the MagMaR (Magnetic Fields in Massive Star-forming Regions) project (Fernández-López et al., 2021; Sanhueza et al., 2021, 2025; Cortés et al., 2021, 2024; Saha et al., 2024; Zapata et al., 2024), which observed the same sample of this study at the same angular resolution than the C-5 configuration (03).